OpenStudy (blackbird02):

Cosine Identities. Find the exact value of cos(x+y), given 0

OpenStudy (campbell_st):

You have 2 different triangles, 1st one contains angle x and has an adjacent side of 4 and hypotenuse of 5. The 2nd triangle has an adjacent side of 12 and hypotenuse of 13. In both triangles you need to find the opposite side using pythagoras'theorem. Then you can find sin(x) and sin(y). When you have the sin ratios you can then use Cos(x+y) = cos(x)cos(y) -sin(x)sin(y) Substitute the exact values forr each ratio to get the answer. Hope it helps

OpenStudy (campbell_st):

Both the angles x and y are 1st quadrant angles since the ratios are positive and the restricted domains

OpenStudy (blackbird02):

so, the opposite side in the first is 3 and on the second its 5, is this correct? @campbell_st

OpenStudy (anonymous):

@blackbird02 yes your right now find the sin(x) and sin(y)

OpenStudy (blackbird02):

@Sambhavvinaykya I'm not sure what sin(x) and sin(y) should be. Is it 4/5 and 2/5 ?

OpenStudy (anonymous):

first draw the two triangles 

OpenStudy (anonymous):

now find sin(x)=opp/hyp and find sin(y)

OpenStudy (anonymous):

your sin(x)=4/5 u were correct but sin(y) is wrong

OpenStudy (blackbird02):

oh, sorry. i meant 5/13 ? I've already solved the exact value. It's 28/65.

OpenStudy (anonymous):

how come